y varies directly as x and inversely as the square of z. y=32 when x=100 and z=5. Find y when x=3 and z=9.
Answer by Boreal(15235)

y=kx/z^2
32=k*100/25=4k
k=8
Use k in the next part
y=8x/z^2
=8*3/81=24/81 or 8/27